# Observation of the quantum equivalence principle for matter-waves

### v3

@@ VII.9 Data analysis for figures 2 and 3
VII.9.1 Figure 2 - Extracting the phase and fitting the data
@@
VII.9.2 Figure 3
In Fig. 3(A), we show the phase derived from the experimental data against three calculations of $\phi(2T)$: the numerical simulation, the analytical approximation and the expected phase in the Gedanken experiment. We calculate the lines of the analytical approximation and the Gedanken experiment using Eq. 4 and Eq. 21 respectively. The line of experimental data was derived by a fit to the data of Fig. 2 (as explained above), and the data points of the numerical simulation were derived as explained in Sec. H. We also plot as a guide to the eye deviations from the $T^{3}$ phase accumulation, by plotting a modified version of the analytical approximation,
$\Delta\phi=\frac{mg^{2}}{3\hbar}\left[\left(\frac{mg^{2}}{3\hbar}\right)^{\frac{\alpha}{3}}T^{3+\alpha}+T^{2}T_{kick}+T\left(T_{kick}^{2}+T_{kick}T_{d}\right)-T_{d}\left(T_{kick}+T_{d}\right)^{2}\right],$ (26)
where we set $\alpha=0.15$ to show 5% deviation from the $T^{3}$ phase accumulation, and the coefficient $\frac{mg^{2}}{3\hbar}^{\frac{\alpha}{3}}$ is added to maintain the correct units. The conditions that define the phase in interest are that the interferometer is closed and the reference arm is stationary. The closing condition is best approximated by the condition of equal areas of current (Eq. 36). We found experimentally, analytically, and numerically, that phase is much more sensitive to small deviations from the equal areas of current than to small deviations from the flatness of the reference arm. Thus to estimate the uncertainty bounds of the phase, we repeat the experiment with an increase (decrease) of the magnetic-pulse current $I_{kick}$ by $0.5\,\%$ (which is our estimated experimental precision on this parameter). We use the result to plot the upper (lower) uncertainty bounds of the phase. The inset shows the rate of phase accumulation $\frac{\partial\phi}{\partial t}$, of the same four cases. To get the derivative of the phase for the numerical simulation, we fit a third-order polynomial on the set points and calculate the analytical derivative of the polynomial. For the experimental data we use the analytical derivative of the polynomial found in Fig. 2. The derivative of the analytical approximation and the Gedanken experiment is calculated by the derivative of Eq. 4 and Eq. 21 respectively. We also plot in the inset the 5% deviation from the $T^{3}$ phase accumulation, this time by plotting the derivative of the modified Gedanken experiment phase accumulation.
$\Delta\phi(2T)=\left(\frac{mg^{2}}{24\hbar}\right)^{\frac{3+\alpha}{3}}(2T)^{3+\alpha}\rightarrow\frac{\partial\phi}{\partial(2T)}=\left(3+\alpha\right)\left(\frac{mg^{2}}{24\hbar}\right)^{\frac{3+\alpha}{3}}(2T)^{2+\alpha},$ (27)
@@ VII.9.2 Figure 3
VII.9.3 Error analysis
@@
$\delta S^{2}=\delta a^{2}\cos^{2}\phi+a^{2}\sin^{2}\phi\delta\phi^{2}.$ (28)
@@
$\left\langle\sin^{2}\phi\right\rangle=\frac{1}{2}[1-\langle\cos(2\phi)\rangle],\quad\left\langle\cos^{2}\phi\right\rangle=\frac{1}{2}[1+\langle\cos(2\phi)\rangle].$ (29)
@@
We obtain
$\delta S^{2}=\frac{1}{2}\left[1+\cos(2\phi)e^{-2\delta\phi^{2}}\right]\delta a^{2}+\frac{1}{2}a^{2}\left[1-\cos(2\phi)e^{-2\delta\phi^{2}}\right]\delta\phi^{2}.$ (31)
To extract the STD of the phase, we need to know the STD of the signal ($\delta S^{2}$), the amplitude and phase of the signal ($a$ and $\phi$), which we get from the fit to the data, and the STD of the amplitude ($\delta a$). If we want to find both amplitude and phase uncertainties from the signal uncertainties for a single point this seems to be impossible because at the right-hand-side we have two variables. So we must either assume a given value of $\delta a$ and then find $\delta\phi$ from the signal uncertainty or we should assume some relation of $\delta a$ and $\delta\phi$ along the different measurement points that may enable determination of both uncertainties from the signal. In our case, we assume that the amplitude uncertainty is negligible compared to the phase uncertainty, so we can find the phase uncertainty from the signal uncertainty. This assumption holds well for $2T>1ms$, where the phase noise is the dominant source of noise in the signal, and is approximately proportional to the phase itself. While for $2T<1\,\rm ms$ this assumption is less accurate, the STD of the signal is very small in this range anyway. As the number of iterations per point is 2-4, resulting in large fluctuations of the STD, we fit the STD to a polynomial of order 3, and use the result of the fit as the STD of the phase in Fig. 3(B). The results of the error analysis are shown in Fig. 10.
Figure 10: Summary of the error analysis. (A) The experimental data, including the raw data of the population vs. time, the fit to the envelope, and the fit to the data. (B) The phase as a function of the holding duration, with error bars indicating the STD of the phase. The STD is calculated using the method of error propagation explained in Sec. I3. (C) The STD of the phase, and fit to a third order polynomial. The result of the fit is used as the STD of the phase in Fig. 3B (D) The relative error of the phase, calculated as the ratio of the STD of the phase to the phase, which has an average value of $1.3\%$.
VII.10 Magnetic gradient pulses
The pulses of the magnetic gradient are formed by running a current through three parallel wires, with alternating current orientation, forming a magnetic quadrupole capable of producing high magnetic gradients with low magnetic fields. The wires are $2\rm\,\mu m$ thick, $40\rm\,\mu m$ wide, and are positioned $100\rm\,\mu m$ apart (center to center), creating a quadrupole $97\rm\,\mu m$ below the surface of the chip (see Fig. 6). The currents are produced by a homemade current source, with high stability and fast response, controlled by an analog signal input. The current source has a maximum slew rate of $30\rm\,{mA/\mu s}$, a maximum current of $1\rm\,A$ and stability of $100\rm\,ppm$. We control the current source by an arbitrary waveform generator, which allows us to create smooth and accurate pulses. To obtain a smooth pulse, we demand that the time derivative $\partial I/\partial t$ is continuous and vanishes at the beginning and end of the pulse. We choose to use a cosine pulse shape, as it fulfills the conditions of a smooth pulse and is analytically solvable when calculating trajectories and phases. The equation describing the current for a pulse starting at $t_{0}$, rise and fall times of $\tau$, separated by a duration $w$ of constant current $A$ is given by
$\mathbf{I}(\mathbf{t}):\begin{cases}0&\text{ for }t<t_{0}\\ \frac{I_{\rm max}}{2}\left(1-\cos\left(\frac{\pi\left(t-t_{0}\right)}{\tau}\right)\right)&\text{ for }t_{0}<t\leq t_{0}+\tau\\ I_{\rm max}&\text{ for }t_{0}+\tau<t\leq t_{0}+\tau+w\\ \frac{I_{\rm max}}{2}\left(1-\cos\left(\frac{\pi\left(t-t_{0}-w\right)}{\tau}\right)\right)&\text{ for }t_{0}+\tau+w<t<t_{0}+2\tau+w\\ 0&\text{ for }t\geq t_{0}+2\tau+w\end{cases}$ (32)
@@ VII.10 Magnetic gradient pulses
The acceleration during the holding pulse is tuned to hold the $|1\rangle$ state against gravity. Measurement of the acceleration of the atoms shows that the required current is $I_{hold}=23.07\pm 1.49\rm\,mA$. To maintain the symmetry of the scheme, the momentum given by the two kick pulses should be equal to the momentum given by the holding pulse, which gives the condition
$a_{kick}=\frac{g}{2\tau_{\rm kick}}T_{h}$ (33)
$a_{\rm hold}=g$ (34)
To improve the response of the current source and reduce the effect of the ambient gradient in the chamber, we set the current between the pulses to a finite value, which we call the idle current, and set its value to $I_{\rm idle}=0.47\,\rm mA$. The idle is on during the entire interferometer sequence. Including the idle current, we get the following requirements for the currents
$\left(I_{\rm kick}-I_{\rm idle}\right)2\tau_{\rm kick}=({I_{\rm hold}-I_{\rm idle}}){T_{h}}$ (35)
$I_{\rm kick}=\frac{I_{\rm hold}-I_{\rm idle}}{2\tau_{\rm kick}}T_{h}+I_{\rm idle}$ (36)
VII.11 Magnetic potential energy of the atoms
@@
$V_{B}(\mathbf{r},t)=m_{F}g_{F}\mu_{B}|B(\mathbf{r},t)|+\alpha_{\rm F,m_{F}}|B(\mathbf{r},t)|^{2}$ (37)
where $g_{F}\mu_{B}/h=0.70\,\rm[MHz/G]$ and $\alpha_{\rm F,m_{F}}$ is a coefficient that depends on $F$ and $m_{F}$, given in table 2. Using Eq. 37, the values in table 2 and the measured transition frequency between $|F=2,m_{F}=2\rangle$ and $|F=2,m_{F}=1\rangle$ of $8.799\,\rm MHz$ we can calculate the magnitude of the bias field $|B|=12.62$ G. To calculate $a_{SOZ}$, we find the magnetic gradient required to hold the $|F=2,m_{F}=1\rangle$ state against gravity, and use it to get the value of $a_{SOZ}=0.10\rm\,m/s^{2}$.
@@ F $m_{F}$ 0 $\pm$1 $\pm$2
Table 2: Values of $\alpha_{\rm F,m_{F}}/h$ in units of $\rm Hz/{G}^{2}$, for calculating the second order Zeeman energy in Eq. 37.
VII.12 Ambient gradient in the chamber
We found a small ambient magnetic gradient in the chamber even when the currents on the atom chip were zero. We measured this ambient magnetic gradient in two methods. The first method is measuring the differential acceleration between the $|1\rangle$ and $|0\rangle$ states in a long time of flight of $20-34\,\rm ms$, after the $\pi/2$ pulse and observing the spatial splitting of the states. We fit the spatial splitting vs time and find $a_{ambient}=0.211\pm 0.005\,\rm m/s^{2}$. The data and the fit appear in Fig. 11. The second method is observing a $T^{3}$ phase accumulation when measuring the pure MW DD scheme (without magnetic gradients from chip wires) for a duration of $4\tau_{DD}=4\,\rm ms$, which can, in the presence of the ambient gradient, create a “self emerging” $T^{3}$ full-loop SGI. In the original $T^{3}$ SGI scheme of [39], the loop is closed by inverting the orientation of the magnetic gradients. In this scheme, the magnetic gradients remain constant, and the $\pi$ pulses invert the spin of the arms to create a closed interferometer. We calculate the prediction for the phase by plugging the durations $(\tau_{DD},2\tau_{DD},\tau_{DD})$ and accelerations $a_{1}=(a_{ambient},0,a_{ambient})+g$, $a_{2}=(0,a_{ambient},0)+g$ of the scheme into Eq. 12 which gives:
$\delta\phi=\frac{m}{\hbar}\tau_{DD}^{3}\left(a_{ambient}^{2}+2ga_{ambient}\right)\,,$ (38)
where $\tau_{DD}=\left(T_{h}+2T_{d}\right)/2$. Using the prediction for the phase we extract $a_{ambient}=0.221\pm 0.005\,\rm m/s^{2}$, which corresponds to a magnetic gradient of $0.68\pm 0.02\,\rm G/cm$. The data and the fit of this measurement appear in Fig. 12
Figure 11: Spatial SG splitting between the state $|1\rangle$ and $|0\rangle$ due to the ambient gradient, vs the time of flight after the $\pi/2$ pulse. Fitting the data to $0.5a_{ambient}TOF^{2}$ gives a value of $a_{ambient}=0.211\pm 0.005\,\rm m/s^{2}$
Figure 12: $T^{3}$ population oscillations in a pure MW DD scheme. By fitting the data to $P(\tau_{DD})=50+50\cos\left[\frac{m}{\hbar}\tau_{DD}^{3}\left(a_{ambient}^{2}+2ga_{ambient}\right)\right]$, we can estimate the acceleration due to the ambient gradient, and get $a_{ambient}=0.221\pm 0.005\,\rm m/s^{2}$.
VII.13 Wave-packet size
@@
A better estimation of wave-packet size is derived using our accurate numerical simulation (Fig. 13). The simulation shows that at half way through the interferometer, the wave-packet size is in the range of $1-1.3\,\rm\mu m$. The maximum splitting achieved is thus 6.5 times the wave-packet width.
Figure 13: Numerical calculation of the wave-packet size and the maximal splitting for different interferometer durations. The wave-packet width is defined by the Gaussian width of the wave packet along the z axis. At the largest holding duration the maximal splitting is $7.5\,\rm\mu m$, and the wave-packet widths are $1\,\rm\mu m$ and $1.3\,\rm\mu m$, for the reference and ballistic wave packets respectively.
@@
VII.14 Einstein equivalence principle for Maxwell waves
The identity of inertial and gravitational mass, usually referred to as the weak equivalence principle [1], led Albert Einstein to a generalized equivalence principle: locally a constant gravitational field is indistinguishable from a constant acceleration.
However, the impossibility of distinguishing between a frame with gravity and the one that is accelerated imposes that the laws of physics must be identical in these two systems. This form of the equivalence principle is usually referred to as the Einstein equivalence principle [2]. Here we neglect the gravitational field of the test mass. When we include the gravitational interaction, for example, for a stellar object, it is called the strong equivalence principle.
In this section we compare and contrast the Kasevich-Chu interferometer (KCI) and the QGI. This analysis brings out most clearly that it is the difference in the beam splitters which leads in the KCI to a confirmation of the Einstein equivalence principle of Maxwell waves [44], whereas in QGI it is a confirmation of the Einstein equivalence principle for de-Broglie waves.
In KCI the beam splitters and mirrors are formed [44] by an optical two-photon transition in a $\Lambda$-system which is detuned with respect to the upper level. Although the frequency difference of the transition between the ground state and the third level is in the microwave regime, the total momentum transfer determined by the wave vector $k$ of the light fields reads $2\hbar k$, since the light fields are counter propagating. We emphasize that in the KCI, the momentum transfer is always $2\hbar k$ independent of the momentum of the scattered atom.
In contrast, in the QGI the beam splitters and the mirrors are realized by two different mechanisms. Indeed, the linear gravitational potential and the magnetic field gradient during the ballistic motion (holding time) act as mirrors where every momentum is, in principle, perfectly inverted. It is this feature which ensures the symmetry between the reference and the ballistic wave packet, as shown in Table I. The beam splitters in the QGI ($\pi/2$ MW pulses) create the coherent superposition between two internal states of the atom without transferring a substantial momentum. It is the magnetic-gradient pulses which launch and stop the ballistic wave packet.
Next we compare and contrast the KCI in the Newtonian and the Einsteinian frame. In the Newtonian frame, the atom in the two paths of the interferometer feels the gravitational acceleration and the phase difference is proportional to the square of the time the atom spends in the interferometer, and linear in acceleration. Since there is no difference in the accelerations in the two paths, the cubic phase in the two paths cancels out. In this description the phase of the electromagnetic fields is constant.
In the Einsteinian frame attached to the atom there is no gravity. However, the frequency of the electromagnetic wave is linearly chirped due to the acceleration of the atom. In this frame the phase shift obviously does not arise from gravity but from the chirping of the laser pulses with the rate given by the gravitational acceleration.
It is important to note that experiments [66] determining the gravitational acceleration in the Newtonian frame vary the chirp rate $a$ and record the phase difference in the KCI as a function of $a$. Since the phase difference is linear in the effective acceleration, it is now given by the difference of the chirp rate and the gravitational acceleration. Hence, when the rate is equal to the gravitational acceleration the phase shift vanishes and is thus independent of the interferometer time.
This technique would not work if the electromagnetic wave would not transform accordingly. Hence, this measurement technique based on the KCI is a manifestation of the Einstein equivalence principle for Maxwell waves.
We conclude by comparing and contrasting this aspect to the QGI. In the KCI, the interaction with the atoms is very sensitive to a resonance condition, due to a significant Doppler shift ${\bf k}\cdot{\bf v}$. Therefore, changing reference frames, which amounts to changing the velocity of the atoms, must be accompanied by an adequate change of the electromagnetic field frequency, and the transformation of the electromagnetic fields of the beam splitters and mirrors plays an important role in the KCI. In contrast, in the QGI, the magnetic-gradient pulses - the Fourier-Transform of which may be considered as an electromagnetic wave having a frequency of about 10 kHz - do not require a resonance condition to interact efficiently with the atoms. Furthermore, the Rabi frequency of the MW pulses responsible for the internal spin splitting and spin inversion is much larger than the Doppler shift. Hence, no transformation of MW fields between the two frames is required in the QGI interferometer.
VII.15 Future experiments with nano-particles

### v4

@@ I Data analysis for figures 2 and 3
I.1 Figure 2 - Extracting the phase and fitting the data
@@
I.2 Figures 3
In Fig. 3(A), we show the phase derived from the experimental data against three calculations of $\phi(2T)$: the numerical simulation, the analytical approximation and the expected phase in the Gedanken experiment. We calculate the lines of the analytical approximation and the Gedanken experiment using Eq. 5 and Eq. S17 respectively. The line of experimental data was derived by a fit to the data of Fig. 2 (as explained above), and the data points of the numerical simulation were derived as explained in Sec. H. We also plot as a guide to the eye deviations from the $T^{3}$ phase accumulation, by plotting a modified version of the analytical approximation,
$\Delta\phi=\frac{mg^{2}}{3\hbar}\left[\left(\frac{mg^{2}}{3\hbar}\right)^{\frac{\alpha}{3}}T^{3+\alpha}+T^{2}T_{kick}+T\left(T_{kick}^{2}+T_{kick}T_{d}\right)-T_{d}\left(T_{kick}+T_{d}\right)^{2}\right],$ (S24)
where we set $\alpha=0.15$ to show 5% deviation from the $T^{3}$ phase accumulation, and the coefficient $\frac{mg^{2}}{3\hbar}^{\frac{\alpha}{3}}$ is added to maintain the correct units. The conditions that define the phase in interest are that the interferometer is closed and the reference arm is stationary. The closing condition is best approximated by the condition of equal areas of current (Eq. 36). We found experimentally, analytically, and numerically, that phase is much more sensitive to small deviations from the equal areas of current than to small deviations from the flatness of the reference arm. Thus to estimate the uncertainty bounds of the phase, we repeat the experiment with an increase (decrease) of the magnetic-pulse current $I_{kick}$ by $0.5\,\%$ (which is our estimated experimental precision on this parameter). We use the result to plot the upper (lower) uncertainty bounds of the phase. The inset shows the rate of phase accumulation $\frac{\partial\phi}{\partial t}$, of the same four cases. To get the derivative of the phase for the numerical simulation, we fit a third-order polynomial on the set points and calculate the analytical derivative of the polynomial. For the experimental data we use the analytical derivative of the polynomial found in Fig. 2. The derivative of the analytical approximation and the Gedanken experiment is calculated by the derivative of Eq. 5 and Eq. S17 respectively. We also plot in the inset the 5% deviation from the $T^{3}$ phase accumulation, this time by plotting the derivative of the modified Gedanken experiment phase accumulation.
$\Delta\phi(2T)=\left(\frac{mg^{2}}{24\hbar}\right)^{\frac{3+\alpha}{3}}(2T)^{3+\alpha}\rightarrow\frac{\partial\phi}{\partial(2T)}=\left(3+\alpha\right)\left(\frac{mg^{2}}{24\hbar}\right)^{\frac{3+\alpha}{3}}(2T)^{2+\alpha},$ (S25)
@@ I.2 Figures 3
Finally, the analysis shows that the predicted prefactor of $\frac{1}{3\hbar}mg^{2}$ is consistent with the data.
I.3 Error analysis
@@
$\delta S^{2}=\delta a^{2}\cos^{2}\phi+a^{2}\sin^{2}\phi\delta\phi^{2}.$ (S26)
@@
$\left\langle\sin^{2}\phi\right\rangle=\frac{1}{2}[1-\langle\cos(2\phi)\rangle],\quad\left\langle\cos^{2}\phi\right\rangle=\frac{1}{2}[1+\langle\cos(2\phi)\rangle].$ (S27)
@@
$\langle\cos(2\phi)\rangle_{\phi=\phi_{0}}=\frac{1}{\sqrt{2\pi}\delta\phi}\int_{-\infty}^{\infty}d\phi e^{-\left(\phi-\phi_{0}\right)^{2}/2\delta\phi^{2}}\cos(2\phi)=\cos\left(2\phi_{0}\right)e^{-2\delta\phi^{2}}$ (S28)
We obtain
$\delta S^{2}=\frac{1}{2}\left[1+\cos(2\phi)e^{-2\delta\phi^{2}}\right]\delta a^{2}+\frac{1}{2}a^{2}\left[1-\cos(2\phi)e^{-2\delta\phi^{2}}\right]\delta\phi^{2}.$ (S29)
To extract the STD of the phase, we need to know the STD of the signal ($\delta S^{2}$), the amplitude and phase of the signal ($a$ and $\phi$), which we get from the fit to the data, and the STD of the amplitude ($\delta a$). If we want to find both amplitude and phase uncertainties from the signal uncertainties for a single point this seems to be impossible because at the right-hand-side we have two variables. So we must either assume a given value of $\delta a$ and then find $\delta\phi$ from the signal uncertainty or we should assume some relation of $\delta a$ and $\delta\phi$ along the different measurement points that may enable determination of both uncertainties from the signal. In our case, we assume that the amplitude uncertainty is negligible compared to the phase uncertainty, so we can find the phase uncertainty from the signal uncertainty. This assumption holds well for $2T>1ms$, where the phase noise is the dominant source of noise in the signal, and is approximately proportional to the phase itself. While for $2T<1\,\rm ms$ this assumption is less accurate, the STD of the signal is very small in this range anyway. As the number of iterations per point is 2-4, resulting in large fluctuations of the STD, we fit the STD to a polynomial of order 3, and use the result of the fit as the STD of the phase in Fig. 3(B). The results of the error analysis are shown in Fig. S5.
Figure S5: Summary of the error analysis. (A) The experimental data, including the raw data of the population vs. time, the fit to the envelope, and the fit to the data. (B) The phase as a function of the holding duration, with error bars indicating the STD of the phase. The STD is calculated using the method of error propagation explained in Sec. I3. (C) The STD of the phase, and fit to a third order polynomial. The result of the fit is used as the STD of the phase in Fig. 3B (D) The relative error of the phase, calculated as the ratio of the STD of the phase to the phase, which has an average value of $1.3\%$.
J Magnetic gradient pulses
The pulses of the magnetic gradient are formed by running a current through three parallel wires, with alternating current orientation, forming a magnetic quadrupole capable of producing high magnetic gradients with low magnetic fields. The wires are $2\rm\,\mu m$ thick, $40\rm\,\mu m$ wide, and are positioned $100\rm\,\mu m$ apart (center to center), creating a quadrupole $97\rm\,\mu m$ below the surface of the chip (see Fig. S1). The currents are produced by a homemade current source, with high stability and fast response, controlled by an analog signal input. The current source has a maximum slew rate of $30\rm\,{mA/\mu s}$, a maximum current of $1\rm\,A$ and stability of $100\rm\,ppm$. We control the current source by an arbitrary waveform generator, which allows us to create smooth and accurate pulses. To obtain a smooth pulse, we demand that the time derivative $\partial I/\partial t$ is continuous and vanishes at the beginning and end of the pulse. We choose to use a cosine pulse shape, as it fulfills the conditions of a smooth pulse and is analytically solvable when calculating trajectories and phases. The equation describing the current for a pulse starting at $t_{0}$, rise and fall times of $\tau$, separated by a duration $w$ of constant current $A$ is given by
@@
To maintain the symmetry of the scheme and satisfy the closing condition, the momentum given by the two kick pulses should be equal to the momentum given by the holding pulse, which gives the condition
$a_{kick}=\frac{a_{\rm hold}}{2\tau_{\rm kick}}T_{h}$ (S31)
The acceleration during the holding pulse $a_{\rm hold}$ is tuned to hold the $|1\rangle$ state against gravity (the levitation condition). To calibrate the holding current $I_{\rm hold}$, so that it satisfies the levitation condition, we measured the acceleration of the atoms in the $|1\rangle$ state as a function of the current in the SG wires. For each value of the current, we measured the position of the atoms along the z-axis as a function of time while being held, and fitted each data set to a parabola. The holding current is adjusted so that the total acceleration is equal to zero, namely, the acceleration due to the holding pulse cancels the acceleration due to gravity and the residual magnetic gradient in the chamber. The measurement is shown in Fig. S6, and yields a value of $I_{\rm hold}=23\,\rm\pm 1.5mA$. As mentioned in the methods section of the main text, while the inaccuracy in satisfying the levitation condition is 6.5%, it corresponds to only 0.42% inaccuracy in the phase accumulation, as the deviation from the levitation condition is introduced squared to the phase.
In order to reduce the effect of the ambient gradient in the chamber, and improve the response of the current source, we set the current between the pulses to a small positive value (and not zero), which we call the idle current. The idle current is set to a value of $I_{\rm idle}=0.47\,\rm mA$, and is on during the entire interferometer sequence. Including the idle current, we get the following requirement for the kick current to satisfy the closing condition of the interferometer:
$\left(I_{\rm kick}-I_{\rm idle}\right)2\tau_{\rm kick}=({I_{\rm hold}-I_{\rm idle}}){T_{h}}$ (S32)
$I_{\rm kick}=\frac{I_{\rm hold}-I_{\rm idle}}{2\tau_{\rm kick}}T_{h}+I_{\rm idle}.$ (S33)
The idle current does not affect the levitation condition and is also included in the numerical simulations.
Figure S6: Calibration of the holding current: (a) Position along the z-axis vs. time for different holding current values. Each data set is fitted with a parabola of the form $z(t)=z_{0}+v_{0}t+\frac{1}{2}at^{2}$, where we set $z_{0}=132$ for all the data sets (note that the chip’s position in this measurement was not calibrated to $z=0$. As we are interested in the shape of trajectories, the constant shift in z is irrelevant). (b) The acceleration of the atoms vs. the holding current. A linear fit predicts a zero acceleration for $I_{\rm hold}=23\,\rm\pm 1.5mA$.
K Magnetic potential energy of the atoms
@@
$V_{B}(\mathbf{r},t)=m_{F}g_{F}\mu_{B}|B(\mathbf{r},t)|+\alpha_{\rm F,m_{F}}|B(\mathbf{r},t)|^{2}$ (S34)
where $g_{F}\mu_{B}/h=0.70\,\rm[MHz/G]$ and $\alpha_{\rm F,m_{F}}$ is a coefficient that depends on $F$ and $m_{F}$, given in table S2. Using Eq. S34, the values in table S2 and the measured transition frequency between $|F=2,m_{F}=2\rangle$ and $|F=2,m_{F}=1\rangle$ of $8.799\,\rm MHz$ we can calculate the magnitude of the bias field $|B|=12.62$ G. To calculate $a_{SOZ}$, we find the magnetic gradient required to hold the $|F=2,m_{F}=1\rangle$ state against gravity, and use it to get the value of $a_{SOZ}=0.10\rm\,m/s^{2}$.
@@ F $m_{F}$ 0 $\pm$1 $\pm$2
Table S2: Values of $\alpha_{\rm F,m_{F}}/h$ in units of $\rm Hz/{G}^{2}$, for calculating the second order Zeeman energy in Eq. S34.
L Ambient gradient in the chamber
We found a small ambient magnetic gradient in the chamber even when the currents on the atom chip were zero. We measured this ambient magnetic gradient in two methods. The first method is measuring the differential acceleration between the $|1\rangle$ and $|0\rangle$ states in a long time of flight of $20-34\,\rm ms$, after the $\pi/2$ pulse and observing the spatial splitting of the states. We fit the spatial splitting vs time and find $a_{ambient}=0.211\pm 0.005\,\rm m/s^{2}$. The data and the fit appear in Fig. S7. The second method is observing a $T^{3}$ phase accumulation when measuring the pure MW DD scheme (without magnetic gradients from chip wires) for a duration of $4\tau_{DD}=4\,\rm ms$, which can, in the presence of the ambient gradient, create a “self emerging” $T^{3}$ full-loop SGI. In the original $T^{3}$ SGI scheme of [35], the loop is closed by inverting the orientation of the magnetic gradients. In this scheme, the magnetic gradients remain constant, and the $\pi$ pulses invert the spin of the arms to create a closed interferometer. We calculate the prediction for the phase by plugging the durations $(\tau_{DD},2\tau_{DD},\tau_{DD})$ and accelerations $a_{1}=(a_{ambient},0,a_{ambient})+g$, $a_{2}=(0,a_{ambient},0)+g$ of the scheme into Eq. S8 which gives:
$\delta\phi=\frac{m}{\hbar}\tau_{DD}^{3}\left(a_{ambient}^{2}+2ga_{ambient}\right)\,$, (S35)
where $\tau_{DD}=\left(T_{h}+2T_{d}\right)/2$. Using the prediction for the phase we extract $a_{ambient}=0.221\pm 0.005\,\rm m/s^{2}$, which corresponds to a magnetic gradient of $0.68\pm 0.02\,\rm G/cm$. The data and the fit of this measurement appear in Fig. S8
Figure S7: Spatial SG splitting between the state $|1\rangle$ and $|0\rangle$ due to the ambient gradient, vs the time of flight after the $\pi/2$ pulse. Fitting the data to $0.5a_{ambient}TOF^{2}$ gives a value of $a_{ambient}=0.211\pm 0.005\,\rm m/s^{2}$
Figure S8: $T^{3}$ population oscillations in a pure MW DD scheme. By fitting the data to $P(\tau_{DD})=50+50\cos\left[\frac{m}{\hbar}\tau_{DD}^{3}\left(a_{ambient}^{2}+2ga_{ambient}\right)\right]$, we can estimate the acceleration due to the ambient gradient, and get $a_{ambient}=0.221\pm 0.005\,\rm m/s^{2}$.
M Wave-packet size
@@
A better estimation of wave-packet size is derived using our accurate numerical simulation (Fig. S9). The simulation shows that at half way through the interferometer, the wave-packet size is in the range of $1-1.3\,\rm\mu m$. The maximum splitting achieved is thus 6.5 times the wave-packet width.
@@
Figure S9: Numerical calculation of the wave-packet size and the maximal splitting for different interferometer durations. The wave-packet width is defined by the Gaussian width of the WP along the z axis. At the largest holding duration the maximal splitting is $7.5\,\rm\mu m$, and the wave-packet widths are $1\,\rm\mu m$ and $1.3\,\rm\mu m$, for the reference and ballistic WPs respectively.
N Future experiments with nano-particles